OpenStudy (anonymous):

(y-7)(y to the 2nd power +7y+49)

OpenStudy (anonymous):

y^3+0y^2+0y-343

OpenStudy (anonymous):

y^3-7^3...?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

y * each part of the second equation+ -7 * each part of the second equation

OpenStudy (anonymous):

y*(y^2+7y+49)= y^3+7y^2+49y -7*(y^2+7y+49)=-7y^2-49y-343

OpenStudy (anonymous):

so do i have to collect like terms to get the answer or what do i have to do

OpenStudy (anonymous):

yup. thats y i put the 0y^2 and 0y in. so you can see it

OpenStudy (anonymous):

the ans is y^3 - 343, making it y^3 - 7^3

OpenStudy (anonymous):

i get the ans after the expanding the equation...

OpenStudy (anonymous):

i mean expression...

OpenStudy (anonymous):

ur good jay. just trying to show how we got the answer

OpenStudy (anonymous):

oh haha ok tutor:)